A galaxy of absolute magnitude M= - 20 is at a distance of 700 kpc. Would it be visible to the unaided eye.
1
Expert's answer
2019-06-03T10:08:48-0400

The apparent magnitude "m" of an object is related to its absolute magnitude "M" by the formula

"m = M + 5 \log_{10} \frac{d}{10\,\text{pc}} \, ,"

where "d" is the distance to the object. Substituting the numbers "M = - 20" and "d = 700\, \text{kpc} = 7 \times 10^5\, \text{pc}" into this formula, we obtain "m \approx 4.2". Since the dimmest objects visible to the unaided eye are assigned "m = 6", our galaxy would be visible.


Answer: yes.